OpenStudy (anonymous):

x/5 - x/2= 3 + 3x/10

OpenStudy (saifoo.khan):

x = -5

OpenStudy (anonymous):

how did you get the answer

OpenStudy (anonymous):

multiply by 10 to simplify 2x - 5x = 30 + 3x then collect like terms -3x - 3x = 30 x= 30/-6 = -5

OpenStudy (anonymous):

thanks